# Lifted Forward Planning in Relational Factored Markov Decision Processes with Concurrent Actions

### v1

@@ A.1 Examples for the Relational Cost Graph
Example 12.
Consider the PRVs $Sick(M)$ and $RemoteWork(M)$ and assume we have a parfactor defined over these two PRVs as well as the PRV $Sick^{\prime}(M)$ for the next state. Then, the relational cost graph consists of two vertices $Sick(M)$ and $RemoteWork(M)$ with an edge between them. Furthermore, we can understand why these two PRVs need to be counted together: A person is (not) sick in the next state dependent on that person being (not) sick and (not) working remote in the current state. We need both values for the correct transition probability.
We now apply the definition of a CRV to Example 12.
Example 13.
The CRV for the clique of $Sick(M)$ and $RemoteWork(M)$ has the structure #$\#[Sick(M),RemoteWork(M)]$ and a histogram for that CRV would consist of the four entries $n_{tt}$ for the number of people being sick and working remote, $n_{tf}$ for the number of people sick but not working remote, $n_{ft}$ for the number of people not sick but working remote, and $n_{ff}$ for the number of people not sick and not working remote.
@@
Example 14.
Suppose we have a parfactor defined over PRVs $Sick(X),Friends(X,Y),Sick(Y)$. The respective CRV has structure #$\#[Sick(X),Friends(X,Y),Sick(Y)]$ and we count for each (joint) assignment to the groundings of the PRVs how often this assignment occurs. Assume that $X$ has a domain with only one person $x$ and $Y$ with three, $y_{1}$ to $y_{3}$. Now further assume that $x$ and $y_{1}$ are sick and $x$ is friends with everyone except $y_{2}$. In other words, $x$ is friends with one sick and one healthy person and not friends with a sick person. Thus, the entries in the histogram for $ttt$, $ttf$ and $tft$ are one, while all other ones are zero.
A.2 Proof of Correct State Representation
@@
The representation in Definition 10 is correct.
@@
Given groundings for $g$, we count the occurrences of each assignment to the PRVs in $\mathbf{B}_{in}$. We split the counts into different CRVs according to Definition 10.
@@
A.3 Calculating Transition Probabilities
In this section, we show how Foreplan calculates the transition probabilities for the constraints in the linear program in Equation 2. Foe each state and action combination, Foreplan generates one constraint. Within this constraint, a sum is taken over all future states. We show how to calculate the required $P(s^{\prime}\mid s,a)$ for given state $s$, action $a$ and next state $s^{\prime}$. We assume that we have one parfactor per PRV in the next state.
@@ A.3 Calculating Transition Probabilities
where $\phi_{i\to true}$ denotes the probability for transitioning from bucket $i$ to $true$ and $\phi_{i\to false}$ for the transition to $false$. Both probabilites can be looked-up in the parfactor. For a parfactor for the output PRV $s_{j}$, the computation is thus
@@
A.4 Example of Removing a Maximum Operator
@@
Note that $e_{2}$ has an empty scope and evaluates therefore to a number. We continue with the second phase, in which ALP translates the elimination sequence into linear program constraints: In the linear program, ALP adds helping variables and constraints to enforce the maxima in the different terms [16]. For each function $e$ with domain $Z$, ALP adds the variables $u_{z}^{e}$ for each assignment $z$ to $Z$. The variable $u_{z}^{e}$ is supposed to yield the value of $e(z)$. For the initial functions $f_{i}$, in our case $f_{1}$, $f_{2}$, $f_{3}$, ALP simply adds $u_{z}^{f_{i}}=f_{i}(z)$ to the constraints of the linear program. Suppose we got the function $e(z)=\max_{x}e_{i}$ when eliminating some variable $x$. ALP then adds the constraints
@@
A.5 More Runtime Theorems for Approximate Foreplan
@@
By definition, the total relational cost graph is a supergraph for all graphs of interest for the runtime complexity:
Theorem 23.
The following graphs are each subgraphs of the total relational cost graph:
  1. 1.
@@
We start with the relational cost graph. By Definition 8, the relational cost graph has a vertex for each PRV and an edge between two PRVs if they share a logvar and occur together in a parfactor, a parameterized local reward function, or a basis function. In particular, all these vertices and edges are introduced in the total relational cost graph too, when we connect two vertices once the corresponding PRVs occur together in a parfactor. We may add more edges than for the relational cost graph as we ignore the share a logvar condition.
@@
Since the induced width is the same as the treewidth minus one, we can give some more bounds:
Theorem 24.
If the total relational cost graph for an rfMDP has bounded treewidth, Approximate Foreplan runs in time polynomial in the number of objects of the rfMDP.
@@
By Theorem 23, the relational cost graph is a subgraph of the total relational cost graph. As a subgraph can have the treewidth of the supergraph at most [22], the treewidth of the relational cost graph is bounded. And if a graph has bounded treewidth, it also has bounded clique number, which is our $w$ [22]. Thus, Theorem 17 follows, leaving only the induced width of each cost network.
By Theorem 23, each cost network is a subgraph of the total relational cost graph. The treewidth of each cost network is bounded by the treewidth of the total relational cost graph. As the induced width equals the treewidth minus one, it is bound, leaving no variable with exponential influence. ∎
A.6 Proof of Approximation Guarantee
@@
A.7 Evaluation
Figure 3 shows the runtime of (Approximate) Foreplan, ALP and XADD Symbolic Value Iteration with a time limit of two hours. The flattening of the runtimes of Foreplan backs our theoretical result in Theorem 15, highlighting that the runtime is indeed not exponential, but rather polynomial, in the number of persons. Moreover, with Approximate Foreplan, we can solve the epidemic example for 164 persons in approximately $2.288$ seconds, which is faster than ALP on 15 persons with approximately $3.045$ seconds and Foreplan on 21 persons with approximately $5.534$ seconds.
Figure 3: Runtime (logscale) of (Approximate) Foreplan, ALP and XADD Symbolic Value Iteration on the epidemic example with a time limit of two hours and a memory limit of 16 GB. Only runs within these limits are shown.
@@ Appendix B Walkthrough of Approximate Foreplan
B.1 Modeling the Small Town
We model the setting in Example 3. The town population consists of three people. We refer to Figure 1 for the transition model. We define the parfactors $f_{1}$ and $f_{2}$ according to Tables 1 and 2.
@@ B.1 Modeling the Small Town
Table 1: Transition probabilities for each person giving the probability of travelling in the next timestep given whether the person is currently travelling and being restricted.
$Sick(M)$ $Epidemic$ $Sick^{\prime}(M)$
@@
Table 2: Transition probabilities for each person giving the probability of being sick in the next timestep given whether the person is currently sick and there is an epidemic.
The mayor can restrict an arbitrary subset of the towns population from travelling. The local reward functions are $R_{1}(Sick(M))$ and $R_{2}(Travel(M))$, each for each person of the towns population. The function $R_{1}(Sick(M))$ evaluates to $1$ if the person is not sick and to $-1$ if the person is sick (c.f. Example 5):
@@
The function $R_{2}(Travel(M))$ evaluates to $2$ if the person is travelling and to 00 otherwise:
@@
B.2 State Space Representation
For using Approximate Foreplan, we first build the relational cost graph to find the state space representation. Afterwards, we can modify the actions to be compatible with the state space representation. The relational cost graph contains two vertices, $Sick(M)$ and $Travel(M)$, because these are the only PRVs in the example. The relational cost graph does not contain any edge, because the two PRVs do not occur together in a parfactor, reward or basis function. Thus, the vertices $Sick(M)$ and $Travel(M)$ each are a clique of size one. Therefore, the state space representation contains two histograms along the value of the propositional random variable $Epidemic$. More formally, the state space representation is $([Sick(M)],[Travel(M)],Epidemic)$ (c.f. Example 8).
For the action $Restrict(M)$, the mayor no longer needs to specify on which person(s) she applies the travel ban(s) on, but has to define how many persons currently (not) travelling receive a travel restriction. Thus, the mayor needs to define the action histogram $[Travel(M),Restrict(M)]$ giving the numbers of people (not) travelling that are (not) restricted from travelling (c.f. Example 9).
B.3 Computing Backprojections
@@
$g_{1}(true,true)$ $=0.6\cdot(-1)+0.4\cdot 1$ =-= - $0.2,$ (31)
@@
$g_{1}(false,true)$ $=0.8\cdot(-1)+0.2\cdot 1$ =-= - $0.6,$ (33)
$g_{1}(false,false)$ $=0.2\cdot(-1)+0.8\cdot 1$ $=$ $0.6$ (34)
@@
$=0.1\cdot h_{2}(t)+0.9\cdot h_{2}(f)=0.1\cdot 2$
$=0.2.$
B.4 Instantiation of Linear Program
@@
$0\geq$ $\max_{x_{1},x_{2},x_{3}}\{3-2\cdot x_{1}+2\cdot x_{2}+w_{0}\cdot(0.9\cdot 1-1)$ (40)
$+w_{1}\cdot(0.9\cdot G_{1}(x_{1},x_{3})-H_{1}(x_{1}))$
$+w_{2}\cdot(0.9\cdot G_{2}^{a}(x_{2})-H_{2}(x_{2}))\},$
@@
$0\geq$ $\max_{x_{1},x_{2},x_{3}}\{3-2x_{1}+2x_{2}-0.1w_{0}$ (41)
$+w_{1}\cdot(x_{1}\cdot 0.9\cdot g_{1}(t,x_{3})-x_{1}\cdot h_{1}(t))$
$+w_{1}\cdot((3-x_{1})\cdot 0.9\cdot g_{1}(f,x_{3})-(3-x_{1})\cdot h_{1}(f))$
$+w_{2}\cdot(x_{2}\cdot 0.9\cdot g_{2}^{f}(t)-x_{2}\cdot h_{2}(t))$
$+w_{2}\cdot((3-x_{2})\cdot 0.9\cdot g_{2}^{f}(f)-(3-x_{2})\cdot h_{2}(f))\}.$
@@ B.4.2 Constraint Generation
B.4.3 Solving the Linear Program

### v2

@@ A.3. Examples for the Relational Cost Graph
[…]
Consider the PRVs $Sick(M)$ and $RemoteWork(M)$ and assume we have a parfactor defined over these two PRVs as well as the PRV $Sick^{\prime}(M)$ for the next state. Then, the relational cost graph consists of two vertices $Sick(M)$ and $RemoteWork(M)$ with an edge between them. Figure 5 shows the graph visually. Furthermore, we can understand why these two PRVs need to be counted together: A person is (not) sick in the next state dependent on that person being (not) sick and (not) working remote in the current state. We need both values for the correct transition probability.
Sick(M)Sick(M)$ReWo(M)$
Figure 5. Visual representation of the relational cost graph in Example A.3. We abbreviate $RemoteWork(M)$ by $ReWo(M)$.
[…]
The relational cost graph consists of two vertices Sick(M) and RemoteWork(M) with an edge between them.
We now apply the definition of a CRV to Example A.3.
[…]
The CRV for the clique consisting of $Sick(M)$ and $RemoteWork(M)$ has the structure #[Sick(M),RemoteWork(M)]\#[Sick(M),RemoteWork(M)] and a histogram for that CRV would consist of the four entries $n_{tt}$ for the number of people being sick and working remote, $n_{tf}$ for the number of people sick but not working remote, $n_{ft}$ for the number of people not sick but working remote, and $n_{ff}$ for the number of people not sick and not working remote.
@@
[…]
Suppose we have a parfactor defined over PRVs $Sick(X),Friends(X,Y),Sick(Y)$. The respective CRV has structure #[Sick(X),Friends(X,Y),Sick(Y)]\#[Sick(X),Friends(X,Y),Sick(Y)] and we count for each (joint) assignment to the groundings of the PRVs how often this assignment occurs. Assume that $X$ has a domain with only one person $x$ and $Y$ with three, $y_{1}$ to $y_{3}$. Now further assume that $x$ and $y_{1}$ are sick and $x$ is friends with everyone except $y_{2}$. In other words, $x$ is friends with one sick and one healthy person and not friends with a sick person. Thus, the entries in the histogram for tttttt, ttfttf and tfttft are one, while all other ones are zero.
A.4. Proof of Correct State Representation
@@
The representation in Definition 4.1 is correct.
@@
Given groundings for $g$, we count the occurrences of each assignment to the PRVs in $\mathbf{B}_{in}$. We split the counts into different CRVs according to Definition 4.1.
@@
A.5. Calculating Transition Probabilities
In this section, we show how Foreplan calculates the transition probabilities for the constraints in the linear program in Equation 2. For each state and action combination, Foreplan generates one constraint. Within this constraint, a sum is taken over all future states. We show how to calculate the required $P(s^{\prime}\mid s,a)$ for given state $s$, action $a$ and next state $s^{\prime}$. We assume that we have one parfactor per PRV in the next state.
@@ A.5. Calculating Transition Probabilities
where $\phi_{i\to true}$ denotes the probability for transitioning from bucket $i$ to truetrue and $\phi_{i\to false}$ for the transition to falsefalse. Both probabilites can be looked-up in the parfactor. For a parfactor for the output PRV $s_{j}$, the computation is thus
@@
A.6. Example of Removing a Maximum Operator
@@
Note that $e_{2}$ has an empty scope and evaluates therefore to a number. We continue with the second phase, in which ALP translates the elimination sequence into linear program constraints: In the linear program, ALP adds helping variables and constraints to enforce the maxima in the different terms approx_factored_mdps. For each function $e$ with domain $Z$, ALP adds the variables $u_{z}^{e}$ for each assignment $z$ to $Z$. The variable $u_{z}^{e}$ is supposed to yield the value of $e(z)$. For the initial functions $f_{i}$, in our case $f_{1}$, $f_{2}$, $f_{3}$, ALP simply adds $u_{z}^{f_{i}}=f_{i}(z)$ to the constraints of the linear program. Suppose we got the function $e(z)=\max_{x}e_{i}$ when eliminating some variable $x$. ALP then adds the constraints
@@
A.7. Proof of Runtime Complexity of Approximate Foreplan
We present the full proof for the runtime complexity of Approximate Foreplan:
Theorem.
Proof.
The runtime of Approximate Foreplan is the runtime of precomputing the backprojections and solving the linear program in Equation 5.
The precomputation involves to calculate the backprojections for each state and action combination. The backprojections are computed as given in Definition 6, which involves another iteration over the state space. Thus, the complexity of precomputing the backprojections is bounded by the state complexity raised to the power of three. However, the size of the state space is, in contrast to Foreplan, no longer exponential in $c$, but only polynomial in $c$: Because of the basis functions, the CRVs of the state representation are not iterated jointly, but independently. For the first part of Approximate Foreplan, we are left with a runtime complexity polynomial in the state space, which is polynomial in the number of objects and $c$, and exponential in $w$.
The second part is solving the linear program in Equation 5. Linear programs can be solved in time polynomial w.r.t the number of variables and constraints 63499. Therefore, we count the number of variables and constraints. Initially, we have one variable per basis function and one constraint per action. Then, the maximum operator in each action constraint is removed. The removal process follows the algorithm by approx_factored_mdps, who have shown their results for arbitrary functions and is therefore applicable. The removal process takes time exponential in the induced width of each cost network approx_factored_mdps; dechter1999bucket. Therefore, the runtime complexity for solving the linear program is polynomial in the number of actions and exponential in the induced width of each cost network.
The overall runtime complexity is (i) polynomial in $c$ and the number of objects, (ii) exponential in the $w$ and the induced width of each cost network. With $w$ bounded, the claim holds. ∎
A.8. More Runtime Theorems for Approximate Foreplan
@@
By definition, the total relational cost graph is a supergraph for all graphs of interest for the runtime complexity: {theorem} The following graphs are each subgraphs of the total relational cost graph:
  1. (1)
@@
We start with the relational cost graph. By Definition 4.1, the relational cost graph has a vertex for each PRV and an edge between two PRVs if they share a logvar and occur together in a parfactor, a parameterized local reward function, or a basis function. In particular, all these vertices and edges are introduced in the total relational cost graph too, when we connect two vertices once the corresponding PRVs occur together in a parfactor. We may add more edges than for the relational cost graph as we ignore the share a logvar condition.
@@
Since the induced width is the same as the treewidth minus one, we can give some more bounds: {theorem} If the total relational cost graph for an rfMDP has bounded treewidth, Approximate Foreplan runs in time polynomial in the number of objects of the rfMDP.
@@
By Theorem A.8, the relational cost graph is a subgraph of the total relational cost graph. As a subgraph can have the treewidth of the supergraph at most db122a323d0740a3b2b1580529f310e7, the treewidth of the relational cost graph is bounded. And if a graph has bounded treewidth, it also has bounded clique number, which is our $w$ db122a323d0740a3b2b1580529f310e7. Thus, Theorem 6 follows, leaving only the induced width of each cost network.
By Theorem A.8, each cost network is a subgraph of the total relational cost graph. The treewidth of each cost network is bounded by the treewidth of the total relational cost graph. As the induced width equals the treewidth minus one, it is bound, leaving no variable with exponential influence. ∎
A.9. Proof of Approximation Guarantee
@@
A.10. Evaluation
Figure 6 shows the runtime of (Approximate) Foreplan, ALP and XADD Symbolic Value Iteration (VI) on the fully-connected SysAdmin approx_factored_mdps example with a time limit of two hours. Both versions of Foreplan clearly surpass ALP and XADD Symbolic VI.
For the BoxWorld boutilier2001symbolic example, we use trains instead of trucks and increase the number of boxes and trains evenly and simultaneously. Figure 7 shows the runtime of (Approximate) Foreplan, ALP and XADD Symbolic VI on the BoxWorld example with a time limit of 15 hours. Foreplan is more than $83.227$ times faster than the exact XADD Symbolic VI, which times out after two boxes and trains. Foreplan is at nine boxes and trains even more than three times faster than ALP. For approximation, Approximate Foreplan is more than $2.057$ times faster than ALP at nine boxes and trains. Moreover, Approximate Foreplan manages to go up to 30 boxes and trains within the same time limit.
Figure 6. Runtime (logscale) of (Approximate) Foreplan, ALP and XADD Symbolic VI on the SysAdmin example with a time limit of two hours. Only runs within these limits are shown. For Approximate Foreplan, we have chosen a bigger step size when no other algorithm was left for the number of computers.
[…]
The figure shows the runtime results of Foreplan, Approximate Foreplan, ALP and XADD Symbolic VI on the SysAdmin example.
Figure 7. Runtime (logscale) of (Approximate) Foreplan, ALP and XADD Symbolic VI on the BoxWorld example with a time limit of 15 hours. Only runs within these limits are shown. For Approximate Foreplan, we have chosen a bigger step size when no other algorithm was left for the number of boxes and trains.
[…]
The figure shows the runtime results of Foreplan, Approximate Foreplan, ALP and XADD Symbolic VI on the BoxWorld example.
The approximation error in the BoxWorld example is more complex. With our chosen basis functions, ALP and Approximate Foreplan both deviate in up to $85$ % of all actions, when tested on up to eight boxes and trains. Still, Approximate Foreplan is not worse than ALP, since both report the same coefficients for the basis functions, empirically and theoretically by Theorem 6. The reason lies in the reward and basis functions: The reward functions award points only for boxes in cities. Because we have chosen the basis functions identical to the reward functions and an additional constant basis function, the basis functions do not take boxes on trains into account, where the boxes may be unloaded in a more profitable city. Therefore, with better basis functions, we should be able to capture better solutions. Moreover, we highlight that Approximate Foreplan returns the same policy as ALP and is more than $30.000$ times faster.
@@ Appendix B Walkthrough of Approximate Foreplan
B.1. Modeling the Small Town
We model the setting in Example 3. The town population consists of three people. We refer to Figure 2 for the transition model. We define the parfactors $f_{1}$ and $f_{2}$ according to Tables 2 and 3. The parfactor $f_{3}$ sets the probability of an epidemic in the next state to $\frac{s+1}{5}$, where $s$ is the number of sick persons in the current state.
Table 2. Transition probabilities for each person giving the probability of travelling in the next timestep given whether the person is currently travelling and being restricted.
@@ B.1. Modeling the Small Town
Table 3. Transition probabilities for each person giving the probability of being sick in the next timestep given whether the person is currently sick and there is an epidemic.
$Sick(M)$ EpidemicEpidemic $Sick^{\prime}(M)$
@@
The mayor can restrict an arbitrary subset of the towns population from travelling. The local reward functions are $R_{1}(Sick(M))$ and $R_{2}(Travel(M))$, each for each person of the towns population. The function $R_{1}(Sick(M))$ evaluates to $1$ if the person is not sick and to 1-1 if the person is sick (c.f. Example 3.2):
@@
The function $R_{2}(Travel(M))$ evaluates to $2$ if the person is travelling and to 0 otherwise:
@@
B.2. State Space Representation
For using Approximate Foreplan, we first build the relational cost graph to find the state space representation. Afterwards, we can modify the actions to be compatible with the state space representation. The relational cost graph, depicted in Figure 3, contains two vertices, $Sick(M)$ and $Travel(M)$, because these are the only PRVs in the example. The relational cost graph does not contain any edge, because the two PRVs do not occur together in a parfactor, reward or basis function. Thus, the vertices $Sick(M)$ and $Travel(M)$ each are a clique of size one. Therefore, the state space representation contains two histograms along the value of the propositional random variable $Epidemic$. More formally, the state space representation is $([Sick(M)],[Travel(M)],Epidemic)$ (c.f. Example 4.1).
For the action $Restrict(M)$, the mayor no longer needs to specify on which person(s) she applies the travel ban(s) on, but has to define how many persons currently (not) travelling receive a travel restriction. Thus, the mayor needs to define the action histogram $[Travel(M),Restrict(M)]$ giving the numbers of people (not) travelling that are (not) restricted from travelling (c.f. Example 4.1).
B.3. Computing Backprojections
@@
$g_{1}(true,true)$ $=0.6\cdot(-1)+0.4\cdot 1$ $=$- $0.2,$ (38)
@@
$g_{1}(false,true)$ $=0.8\cdot(-1)+0.2\cdot 1$ $=$- $0.6,$ (40)
$g_{1}(false,false)$ $=0.2\cdot(-1)+0.8\cdot 1$ $=$ $0.6.$ (41)
@@
[…]
B.4. Instantiation of Linear Program
@@
[…]
@@
[…]
@@ B.4.2. Constraint Generation
B.4.3. Solving the Linear Program